The following company information is available for March. The direct materials price variance is:

Direct materials purchased and used 4,100 feet @ $93 per foot
Standard costs for direct materials for March production 4,200 feet @ $91 per foot

Multiple Choice

  • $8,200 unfavorable.

  • $8,200 favorable.

  • $8,400 unfavorable.

  • $8,400 favorable.

  • $900 favorable
Answer #1

Materials Price Variance

= (SP-AP) *AQ

= (91-93)*4100

= - 2*4100

= 8200 Unfavorable